# Phonon thermal Hall as a lattice Aharonov-Bohm effect

### v1

@@ Abstract
In a growing list of insulators, experiments find a misalignment between the heat flux and the thermal gradient vectors induced by magnetic field. This phenomenon, known as the phonon thermal Hall effect, implies energy flow without entropy production along the orientation perpendicular to the temperature gradient. Experiments find that the thermal Hall angle is maximal at the temperature at which the longitudinal thermal conductivity peaks. At this temperature, $T_{max}$, Normal phonon-phonon collisions (which do not produce entropy) dominate Umklapp and boundary scattering events (which do). In the presence of a magnetic field, Born-Oppenheimer approximated molecular wave functions are known to acquire a geometric [Berry] phase. I will argue here that the survival of this phase in a crystal implies a complex amplitude for transverse phonons. This modifies three-phonon interference patterns, twisting the quasi-momentum of the outgoing phonon. The rough amplitude of the thermal Hall angle expected in this picture is set by the wavelength, $\lambda_{max}$, and the crest displacement amplitude, $u_{m}$, of transverse acoustic phonons at $T_{max}$. Combined with the interatomic distance, $a$ and the magnetic length, $\ell_{B}$, it yields: $\Theta_{H}\approx\lambda_{max}^{2}u_{m}^{2}a^{-2}\ell_{B}^{-2}$. This is surprisingly close to what has been experimentally found in black phosphorus, germanium and silicon.
@@ 3 Anharmonicity and phonon-phonon collisions
Absorption and emission events are distinct only in presence of an arrow of time. If instead of beginning with $q_{1}$ and ending with $q_{3}$ one begins with the latter and ends with the former, an emission event becomes an absorption event and vice versa. Such events do not produce entropy. However, there are ph-ph scattering events, which do produce entropy (See Fig. 3b). If $\overrightarrow{q_{3}}$ gets out of the Brillouin zone, the scattering event is known as Umklapp. Since a lattice cannot host a sound wave with a wavelength shorter than twice the interatomic distance, $a$, it is impossible to have $|\overrightarrow{q_{3}}|>\frac{\pi}{a}$. When this is the case, a fraction of the quasi-momentum is absorbed by the crystal. Note that in this case, there is no way to start with the forbidden final wave-vector.
@@ 4 From the Born-Oppenheimer approximation to the geometric phase of phonons in a magnetic field
In 1988, Schmelcher, Cederbaum and Meyer [35, 36] examined the fate of the Born-Oppenheimer approximation in the presence of the magnetic field. They found that the simple extension of zero-field Born-Oppenheimer adiabatic approximation with no modification would make nuclei appear as “naked" charges, i.e., with no screening of the magnetic field by the electrons. This led them to to distinguish between the diagonal and off-diagonal terms of the non-adiabatic coupling and formulate what they called the ‘screened Born-Oppenheimer approximation’ [35]. In 1992, Mead [38] argued that this approach originally identified as introducing “diagonal non-adiabatic correction" [35], and motivated by keeping gauge consistency, is equivalent to the addition of a geometric vector potential due to the finite magnetic field. The fate of this phase in molecules subject to strong magnetic field has been studied numerically in recent years [76, 77]. On the other hand, what happens in a crystalline lattice has been barely explored (see [43] for an exception).
Figure 4: a) The hydrogen molecule in presence of a magnetic field. The molecular wavefunction has a phase which is given by the vector product of the magnetic field and the coordinate of each electron with respect to the center of mass of the molecule. This phase arises because the magnetic flux seen by the electron cloud is much larger than the magnetic flux seen by point-like nuclei. b) A transverse acoustic phonon is associated with two length scales, the wavelength $\lambda$ and the maximum atomic displacement, $\delta u_{max}$. It distorts the total amount of the magnetic flux seen by electrons and this distortion is proportional to the product of these two length scales. c) Spheres with different colors represent silicon atoms in their tetrahedral environment. Arrows indicate displacements of the center of masses associated with the two lattice waves encountering each other. These two phonons have a geometric phase shift.
@@
Here $\overrightarrow{r_{l}}$ represent the coordinates of the electron indexed $l$ with respect to the center of mass of the molecule. The latter’s coordinates are represented by $\overrightarrow{R_{S}}$ (See Figure 4a). The ultimate source of this phase factor is the difference in the spatial location of the positive charge of the nuclei and the negative charge of the electron cloud. The Lorentz force exerted by the magnetic field on the electrons and the nuclei does not compensate each other. In comparison to Coulomb interaction, this is a very small correction in laboratory magnetic fields [76, 77]. However, the difference between the magnetic flux seen by the electron cloud and by point-like nuclei generates a qualitative difference with the situation at zero magnetic field. Now, the molecular wave-function cannot be described by a real function in space and time and has a geometric phase.
Does this phase survive in a crystal? If yes, what are the consequences?
Let us consider a transverse phonon with a wavelength of $\lambda$, with atomic displacement of $\delta u_{m}$ at its crest (Fig.4b)). Such a lattice wave distorts the spatial distribution of charge. The electronic cloud of covalent bonds will vibrate. Let us call $Q=q_{e}e$ the ‘shuffled’ amount of electric charge ($q_{e}$ is a real number and $e$ is the fundamental electric charge). What is the change in the geometric phase associated with such a wave? A reasonable guess (inspired by equation 15) is to equate this phase with the change in the [Aharonov-Bohm] magnetic flux times the relevant electric charge divided by the Planck constant and the flux is the product of the magnetic field and the relevant area. Therefore, neglecting numerical prefactors of the order of unity, the shift introduced by a phonon is:
@@
Here, $\ell_{B}=\sqrt{\frac{\hbar}{eB}}$ is the magnetic length. In our temperature and field range of interest, $\ell_{B}^{2}\gg\lambda_{ph}\delta u_{m}$ and therefore, $\delta\phi_{B}\ll 1$. However, such a small phase shift can affect the interference between two colliding phonons (Fig.4c ). In the same magnetic field, two phonons with the same wave-vector but with different polarizations do not sweep the same change in the magnetic flux. Their $\delta\phi_{B}$ is different.
@@ 6 The rough amplitude
A rigorous answer to this question is a task for ab initio studies starting from the phonon spectrum in each solid. In 2019, Saito and collaborators [43] proposed that in a nonmagnetic insulator with ordinary phonons, such as silicon, Berry phase can give rise to a detectable thermal Hall effect. However, their expected thermal hall conductivity was $10^{-6}WK^{-1}m^{-1}$. As seen in table 1, the experimentally observed thermal Hall conductivity in silicon [15] is 7 orders of magnitude larger! Anharmonicity, which is the central player in our scenario, was absent in their model [43]. The amplitude of the thermal Hall conductivity correlates with the maximum longitudinal thermal conductivity and the latter is known to depend on the sample size [62]. The maximum thermal conductivity is not an intrinsic property. It is higher [80] in silicon crystals larger than the one in which the thermal Hall conductivity was measured [15].
In our picture, it is the Hall angle, (the ratio of longitudinal to transverse thermal conductivities at Tmax which is intrinsic (See table 1). One simply expects to see $\Theta\approx 2\delta\phi_{B}$.
@@
Solid Mass (a.u.) a(Å) $v_{t}$(km/s) $\lambda_{t}$ (Å) $\delta u_{m}$(Å) $q_{e}$ $q_{e}\lambda_{t}\Lambda_{M}\ell_{B}^{-2}(10^{-4}T^{-1})$
@@
Table 2: Atomic masses and length scales of silicon, germanium and phosphorous. $a=(V_{c}/n_{c})^{1/3}$ is the average distance per atoms (with $V_{c}$ and $n_{c}$ are the volume of the primitive cell and its number of atoms). $v_{t}$ is the transverse sound velocity used to extract $\lambda_{t}$ at T=$T_{max}$. The two values for Black P represent the lowest and the largest known [62]. $\Lambda_{M}$ is the de Broglie thermal length at T=$T_{max}$ extracted from the atomic mass. The last column is the the rough amplitude of the expected Hall angle in a given magnetic field to be compared with the last column of table 1 (See text).
To quantify $\delta\phi_{B}$, we need to know the three quantities entering in equation 16. For a transverse phonon with linear dispersion with a wavevector of $q_{t}=\frac{2\pi}{\lambda_{t}}$ and a velocity of $v_{t}=\omega_{t}/q_{t}$, the wavelength at a temperature of $T=T_{max}$ is:
$\lambda_{ph}\approx\frac{hv_{t}}{k_{B}T_{max}}$ (19)
The atomic displacement at the crest of a lattice wave [82] can be estimated by quantifying its link to the energy of the phonon mode: $M\delta u_{m}^{2}\omega_{t}^{2}=k_{B}T_{max}$, where $M$ is the atomic mass. Combined with $\hbar\omega_{t}(T_{max})=k_{B}T_{max}$, it becomes:
$\delta u_{m}\approx\frac{\hbar}{\sqrt{Mk_{B}T_{max}}}$ (20)
This makes $\delta u_{m}$ proportional to the de Broglie thermal length of the vibrating atom.
Among the three parameters of equation 16, estimating $q_{e}$ is the least straightforward. Let us call $a$ the interatomic distance. In the simplest approach, the number of atoms perturbed by a transverse phonon is $\approx\lambda_{t}/a$. For each of these atoms, $\delta u_{m}/a$ represents the fraction of electron cloud concerned by atomic displacement. This leads to:
$q_{e}\approx\frac{\delta u_{m}\lambda_{t}}{a^{2}}$ (21)
@@ 7 Concluding remarks
In [44], Li, Guo, Zhu and the present author proposed that electric dipolar waves traveling with phonons can generate a thermal Hall effect. While the narrative thread of the present paper is different, the core idea is similar. After all, what turns the hydrogen atom to an electric dipole is the existence of an electronic cloud and a point-like nucleus. For exactly the same reason, the hydrogen atom has a geometric phase in a magnetic field. The present scenario, which invokes phonon-phonon interference as a consequence of a complex phase, can provide an account of the temperature dependence and the order of magnitude of the observed signal.
@@
To this very pertinent judgment, one may add that it is also remarkable that their consequences for phonons have not given much thought. If the phase of the wavefunction matters for molecules, why should it be neglected for a sound wave in a crystal ? The absence becomes even more surprising, given that Brown [83] had explicitly found an additional phase factor between the zero-field and finite field translation operators for Bloch electrons in a uniform magnetic field. There is no trace of the implications of this phase factor for waves in insulators.

### v2

@@ Abstract
In a growing list of insulators, experiments find a misalignment between the heat flux and the thermal gradient vectors induced by magnetic field. This phenomenon, known as the phonon thermal Hall effect, implies energy flow without entropy production along the orientation perpendicular to the temperature gradient. Experiments find that the thermal Hall angle is maximal at the temperature at which the longitudinal thermal conductivity peaks. At this temperature, $T_{max}$, Normal phonon-phonon collisions (which do not produce entropy) dominate Umklapp and boundary scattering events (which do). In the presence of a magnetic field, Born-Oppenheimer approximated molecular wave functions are known to acquire a geometric [Berry] phase. Its origin is the difference in the spatial distributions of the positive charge of the nuclei and the negative charge of the electron cloud. The magnetic flux is negligible for the point-like nuclei, but not for the electron cloud. Therefore, the molecular wave-function becomes a complex number in the presence of the magnetic field. In a crystal, this implies a complex amplitude for transverse acoustic phonons, modifying three-phonon interference patterns, and twisting the quasi-momentum of the outgoing phonon. The rough amplitude of the thermal Hall angle expected in this picture is set by the wavelength, $\lambda_{ph}$, and the crest displacement amplitude, $u_{m}$, of transverse acoustic phonons at $T_{max}$. The derived expression is surprisingly close to what has been experimentally found in black phosphorus, germanium and silicon.
@@ 3 Anharmonicity and phonon-phonon collisions
Absorption and emission events are distinct only in the presence of an arrow of time. If instead of beginning with $q_{1}$ and ending with $q_{3}$ one begins with the latter and ends with the former, an emission event becomes an absorption event and vice versa. Such events do not produce entropy. However, there are ph-ph scattering events, which do produce entropy (See Fig. 3b). If $\overrightarrow{q_{3}}$ gets out of the Brillouin zone, the scattering event is known as Umklapp. Since a lattice cannot host a sound wave with a wavelength shorter than twice the interatomic distance, $a$, it is impossible to have $|\overrightarrow{q_{3}}|>\frac{\pi}{a}$. When this is the case, a fraction of the quasi-momentum is absorbed by the crystal. Note that in this case, there is no way to start with the forbidden final wave-vector.
@@ 4 From the Born-Oppenheimer approximation to the geometric phase of phonons in a magnetic field
In 1988, Schmelcher, Cederbaum and Meyer [35, 36] examined the fate of the Born-Oppenheimer approximation in the presence of the magnetic field. They found that the simple extension of zero-field Born-Oppenheimer adiabatic approximation with no modification would make nuclei appear as “naked" charges, i.e., with no screening of the magnetic field by the electrons. This led them to to distinguish between the diagonal and off-diagonal terms of the non-adiabatic coupling and formulate what they called the ‘screened Born-Oppenheimer approximation’ [35]. In 1992, Mead [38] argued that this approach originally identified as introducing “diagonal non-adiabatic correction" [35], and motivated by keeping gauge consistency, is equivalent to the addition of a geometric vector potential due to the finite magnetic field. The fate of this phase in molecules subject to strong magnetic field has been studied numerically in recent years [76, 77]. On the other hand, what happens in a crystalline lattice has beend barely explored (see [43] for an exception).
Figure 4: a) In the presence of magnetic field, an atom (top), a molecule (middle) and a chain of atoms (bottom) have a gauge-dependent phase. Expressed in the symmetric gauge, it is given by the vector product of the magnetic field and the coordinate of each electron with respect to the center of mass of the molecule. The ultimate origin of this phase is the fact that the magnetic flux swept by the electron cloud and by the point-like nuclei do not cancel out. b) A transverse acoustic phonon distorts the total amount of the magnetic flux swept by electrons. This distortion depends on two length scales, the wavelength $\lambda$ and the maximum atomic displacement, $\delta u_{max}$. These two length scales determine the change in the swept magnetic flux divided by the quantum of magnetic flux and set the phonon phase.
@@
Here $\overrightarrow{r_{l}}$ represent the coordinates of the electron indexed $l$ with respect to the center of mass of the molecule. The latter’s coordinates are represented by $\overrightarrow{R_{S}}$.
This expression invites several comments. Its first remarkable feature is that it does not depend on the ratio of the electron-to-nucleus mass ratio. Note also that this expression is gauge-dependent and it remains to be seen that it has any observable consequences. Moreover, as illustrated in Fig. 4a, such a gauge-dependent phase factor is present in an atomic wave-function, or in the wave-function of a chain of atoms. In each case, one needs to shift the coordinate of the center-of-mass and set the amplitude of the pseudomementum as the sum of atomic pseudomomenta [35]. Finally, let us note that this phase factor of Equation 15 is the ratio of two areas: ($\sum_{l}{r_{l}}.R_{S}$ and $\ell_{B}^{2}$, with $\ell_{B}=\sqrt{\frac{\hbar}{eB}}$ being the magnetic length.)
The ultimate source of this phase factor is the difference in the spatial distributions of the positive charge of the nuclei and the negative charge of the electron cloud. The magnetic flux, which is gauge-invariant, is negligible for the point-like nuclei, but not for the electron cloud. The absence of compensation makes the molecular wave-function inherently complex in the presence of the magnetic field.
In order to see the consequences of this field-induced phase in a crystal, let us consider how a transverse acoustic phonon can change the magnetic flux swept by the electronic cloud of an insulating solid. The two relevant length scales are the phonon wavelength $\lambda_{ph}$, and the atomic displacement of $\delta u_{m}$ at the wave crest (Fig.4b)). Their product defines an area in which atoms are displaced from their equilibrium position. The geometric gauge-independent phase of such a phonon, indicated by our interpretation of equation 15, is this product divided by $\ell_{B}^{2}$. Equation 15 contains a sum of all electron coordinates too. Therefore, another relevant quantity is the total ‘shuffled’ amount of electric charge. Let us call it $Q=q_{e}e$, where $q_{e}$ is a real number and $e$ is the fundamental electric charge. This leads us to :
@@
Let us now see how large $\delta\phi_{B}$ can be. When B $\simeq 10$ T and $T\sim 10K$, $\ell_{B}$ and $\lambda_{ph}$ are of the same order of magnitude (in the range of hundreds of angstroms) and $\delta u_{m}$ is two to three orders of magnitude smaller than both. As we will see in section 6, the dimensionless $q_{e}$ is of the order of unity. Therefore, in our range of interest the phonon phase is of the order of a few miliradians, comparable to the phase shift of an optical photon in a typical Kerr experiment on a magnetic material [78].
This small phase shift can affect the interference between two colliding phonons (Fig.4c). In the same magnetic field, two phonons with the same wave-vector but with different polarizations do not sweep an identical magnetic flux and therefore their $\delta\phi_{B}$ is different.
@@ 6 The rough amplitude
A rigorous answer to this question is a task for ab initio studies starting from the phonon spectrum in each solid. In 2019, Saito and collaborators [43] proposed that in a nonmagnetic insulator with ordinary phonons, such as silicon, Berry phase can give rise to a detectable thermal Hall effect. However, their expected thermal hall conductivity was $10^{-6}WK^{-1}m^{-1}$. As seen in table 1, the experimentally observed thermal Hall conductivity in silicon [15] is 7 orders of magnitude larger! Anharmonicity, the central player in our scenario, was not mentioned there [43].
The amplitude of the thermal Hall conductivity correlates with the maximum longitudinal thermal conductivity. The latter is known to depend on the sample size [62]. The maximum thermal conductivity is not an intrinsic property. In silicon crystals [81] larger than the one in which the thermal Hall conductivity was measured [15], longitudinal thermal conductivity is several times larger.
In the picture drawn here, the Hall angle, the ratio of longitudinal to transverse thermal conductivities at Tmax (see table 1), is the central quantity to put under scrutiny. If the field-induced misalignment between the heat current and the temperature gradient is due to interference between phonons, then one expects it to be $\approx 2\delta\phi_{B}$.
@@
Solid Mass a $v_{t}$ $\lambda_{ph}(T_{max})$ $\delta u_{m}(T_{max})$ $q_{e}(T_{max})$ $\delta\Phi_{B}(T_{max})$
(a.u.) (Å) (km/s) (Å) (Å) $(10^{-4}T^{-1})$
@@
Table 2: Atomic masses and length scales of silicon, germanium and phosphorous. $a=(V_{c}/n_{c})^{1/3}$ is the average distance per atoms (with $V_{c}$ and $n_{c}$ are the volume of the primitive cell and its number of atoms). $v_{t}$ is the transverse sound velocity used to extract the transverse phonon wavelength at T=$T_{max}$. The two $v_{t}$ values for Black P represent the lowest and the largest known [62]. $\delta u_{m}$ is the crest atomic displacement at T=$T_{max}$. The last column is the the rough amplitude of the phonon phase shift at $T_{max}$, to be compared with the last column of table 1.
Let us now quantify $\delta\phi_{B}$ in the three elemental insulators listed in table 1. Let us examine, $\lambda_{ph}$, $\delta u_{m}$ and $q_{e}$, the three quantities to be injected in equation 16.
The most familiar quantity is the wavelength of an acoustic phonon. Consider a linearly dispersive transverse phonon with a wavevector of $q_{t}=\frac{2\pi}{\lambda_{t}}$ and a velocity of $v_{t}=\omega_{t}/q_{t}$. Its wavevector is simply proportional to thermal energy and its wavelength proportional to the inverse of temperature. At, $T=T_{max}$, this wavelength becomes equal to :
$\lambda_{ph}(T_{max}\approx\frac{hv_{t}}{k_{B}T_{max}}$ (19)
Note that at T=0, $q_{t}$ vanishes and the wavelength of an acoustic phonon becomes infinite. This of course never happens in a solid of finite size. However, in realistic experimental set-ups, we are far from this limit. Given that sound velocity in all known solids is between 1 to 20 km/s, the wavelength of thermally excited phonons is about in the range of $\sim 10$ nm at 1 K and $\sim 1\mu$ m at 10 mK. The samples studied in these experiments are much larger.
The amplitude of $\delta u_{m}$, the atomic displacement at the crest of a lattice wave can be estimated by its link to the energy of the thermally excited phonon mode: $M\delta u_{m}^{2}\omega_{t}^{2}=k_{B}T_{max}$, where $M$ is the atomic mass [83]. Combined with $\hbar\omega_{t}(T_{max})=k_{B}T_{max}$, it becomes:
$\delta u_{m}(T_{max})\approx\frac{\hbar}{\sqrt{Mk_{B}T_{max}}}$ (20)
Thus, $\delta u_{m}$ proportional to the de Broglie thermal length of the vibrating atom. Both diverge at T=0. However, this divergence is even weaker than the zero-temperature divergence of $\lambda_{ph}(T)$. In our temperature range of interest, $\delta u_{m}$ remains shorter than the interatomic distance, which we call, $a$ (See table 2).
Among the three parameters of equation 16, estimating $q_{e}$ is the least straightforward. One one hand, an acoustic phonon, which has a wavelength much longer than the interatomic distance, will displace many electrons from their equilibrium position. However, the induced change in the magnetic flux swept by the electronic cloud is a tiny fraction of the electron areal density. The dimensionless $q_{e}$ is set by the product of these two numbers, one much larger than and the other much smaller than unity. The number of atoms perturbed by a transverse phonon with a wavelength of $\lambda_{t}$ is $\approx\lambda_{t}/a$. For each of these atoms, $\delta u_{m}/a$ represents the elongation of electronic bonds between atoms induced by the sound wave. Assuming that each atom shares one electron with its neighbor, this leads to
$q_{e}\approx\frac{\delta u_{m}\lambda_{ph}}{a^{2}}$ (21)
We saw above that $\delta u_{m}$ and $\lambda_{ph}$ both diverge at zero temperature. Therefore, this will also be true of their product and $q_{e}$. However, even at T=1 K, $\frac{\delta u_{m}(T)}{a}<1$ and therefore $q_{e}$, remains a two digit number. Note that the decrease in the population of thermally excited phonons decreases as $\propto T^{3}$ , much faster than the $T^{3/2}$ enhancement of $q_{e}$ At 1 K, phonon-phonon interference is rare and almost all phonons travel ballistically without meeting other phonons.
@@ 7 Concluding remarks
In [44], Li, Guo, Zhu and the present author proposed that electric dipolar waves traveling with phonons can generate a thermal Hall effect. While the narrative thread of the present paper is different, the core idea in both is based on the combination of an electronic cloud and a point-like nucleus. The present scenario, which invokes phonon-phonon interference as a consequence of a complex phase, can provide an account of the temperature dependence and the order of magnitude of the observed signal.
@@
To this very pertinent judgment, one may add that it is also remarkable that their consequences for phonons have not given much thought. If the phase of the wavefunction matters for molecules, why should it be neglected for a sound wave in a crystal ? The absence becomes even more surprising, given that Brown [84] had explicitly found an additional phase factor between the zero-field and finite field translation operators for Bloch electrons in a uniform magnetic field, . There is no trace of the implications of this phase factor for waves in insulators.